Becky and Rita are 648 miles apart. Becky walks toward Rita at the rate of 2.25 m/s and Rita runs towards Becky. What is Rita's average speed if she reaches Becky in 1.6 minutes?
Answer by stanbon(75887)

Becky and Rita are 648 miles apart. Becky walks toward Rita at the rate of 2.25 m/s and Rita runs towards Becky. What is Rita's average speed if she reaches Becky in 1.6 minutes?
---------------
Comment: I don't know if your 2.25 m/s is meters/second or 2.25 miles/second
----------------
Becky DATA:
Rate = 2.25 miles/sec ; distance = x miles ; time = d/r = x/2.25 seconds
---------------
Rita DATA:
Time = 96 seconds ; distance = 648-x miles ; rate = d/t = (648-x)/96 miles/sec.
---------------------------------
EQUATION:
Becky time = Rita time
x/2.25 = 96
x = 216 miles (Becky's distance)
Rita's speed = (648-216)/96 = 432/96 = 4.5 miles/second
===================
Cheers,
Stan H.